# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ Abstract.
The distribution of the modulus of the extreme eigenvalues is investigated for the complex Ginibre and complex induced Ginibre ensembles in the limit of large dimensions of random matrices. The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution. The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes. In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively. The limiting left tail distribution of the minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero. This phenomenon is also verified for the right tail distribution of this minimum.
@@ 1. Introduction
The attractiveness of the Random Matrix Theory (RMT) lies in the possibility of using it as a means to model problems in high dimensions and perform related calculations analytically. This is mainly due to the invariance property of the probability distribution of certain matrix ensembles. As stated in [19], the statistical properties of the spectrum of random matrix ensembles are independent of the nature of the probability distribution that defines these ensembles in the limit of large sizes of these matrices. They only depend on the invariance of these distributions. Matrix ensembles that are characterised by invariant probability distribution are those from classical compact matrix groups studied in references [30] and [17]. The complex Ginibre ensemble is one of them and is a special case of the Ginibre-Girko ensemble with maximal non-Hermiticity as presented in [8] and [2].
The complex Ginibre ensemble was first defined as a mathematical concept to model phenomena from the physics of particles. Statistical properties of eigenvalues of matrices from this ensemble are studied to provide an understanding by analogy of the dynamics of nuclei. More precisely, as stated in [10], the distribution of eigenvalues of a complex Ginibre matrix is comparable to that of the distribution of the positions of charges of a two-dimensional Coulomb gas in a harmonic oscillator potential, at a specific temperature corresponding to the Dyson index $\beta=2$.
The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$ whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density
@@
where $(x_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$ and $(y_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$. The real and imaginary parts of the entries denoted $x_{ij}$ and $y_{ij}$ are independent random variables, each following a real Gaussian distribution $\mathcal{N}_{\mathbb{R}}(0,\frac{1}{2})$. Like the real and the symplectic Ginibre ensembles, this ensemble was introduced by J. Ginibre [10]. The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ is the joint probability density function of the entries $J_{ij}$ defined as
@@
and $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [15].
The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in close form as
@@
where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$. The term $C_{N}$ is the normalisation factor defined as
@@
A topic which has been widely investigated in RMT is the limiting distribution of the eigenvalues of random matrix ensembles as $N$ goes to infinity. The notion of universality in this field has found relevance in establishing analytical expressions of the limiting distribution of random matrix eigenvalues. It is a mathematical concept asserting that the limiting distribution of the eigenvalues should not depend on the particular distribution of the random matrix entries [23]. This has led to several important results among which is referenced the circular law conjecture for non-Hermitian random matrix ensembles (cf. [11], [23], [3]).
Studies of the limiting distribution of the largest and smallest eigenvalues for Gaussian ensembles are presented in [27, 28, 29]. The distributions of eigenvalues (and their spacings) have been investigated with the computation of gap probabilities with respect to radial ordering for non-Hermitian random matrices and their chiral counterparts [2]. The statistical properties of extreme eigenvalue moduli of matrices from non-Hermitian ensembles have also been studied in the literature such as the limiting distribution of the spectral radius for the complex Ginibre ensemble [20, 21] and the real Ginibre ensemble in [22].
Section 2 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble. The main results are derived for the spectral radius and the minimum modulus. The limiting left and right tail distributions of the minimum modulus is also studied for large size of matrices in Section 2.1. Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.2. The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in [16], and is extended in [14].
Employing the method introduced in [20] from which limit theorems are derived at the edge of the spectrum for the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [5] are presented in Section 3. This random matrix ensemble might find relevance for questions raised in different fields of physics. It would also define an ideal modelling framework for the expansion of the research related to non-Hermitian matrices. The complex induced Ginibre ensemble is a special case of the Feinberg-Zee ensemble and a generalisation of the complex Ginibre ensemble. The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is
@@
The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V(y)=-\frac{\beta}{2}(y-L\log(y))$ where $L$ is a non-negative parameter called the rectangularity index.
@@
$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$ (6)
where $\beta=2$ for the complex induced Ginibre ensemble and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$. The factor $\Gamma^{(\beta)}$ is the normalisation constant
@@
In the limit of large matrices dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble thus referenced as the circular law and presented in the work from V. L. Girko [11] and the paper from T. Tao, V. Vu and M. Krishnapur [23]. The joint probability density function of the eigenvalues of complex induced Ginibre matrices is
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$ (8)
@@
It is understood from the definition of the joint probability density function of the eigenvalues that there is a repulsion of the eigenvalues from the origin of the complex plane due to the term $\det(GG^{\dagger})^{L}=\prod_{j=1}^{N}|\lambda|^{2L}$ appearing in the formula. This phenomenon is explained from a mathematical perspective with the formulation of the mean eigenvalue density, i.e., $\langle\rho_{N}(\lambda)\rangle=R_{1}(\lambda)$ where $R_{1}(\lambda)$ is the one-point correlation function of the eigenvalues. More precisely taking its limit as $N$ goes to infinity and for a rectangularity index $L$, the distribution of the eigenvalues is uniform and supported by a ring centred at the origin of the complex plane with a outer circle of radius $r_{out}=\sqrt{L+N}$ and an inner circle of radius $r_{in}=\sqrt{L}$. For small rectangularity indexes $L$ (close to zero), the density of the eigenvalues is uniform over the disk of radius $\sqrt{N}$ with a mean eigenvalue density converging to the mean eigenvalue density of the complex Ginibre ensemble.
@@
In the present paper, an investigation is performed for the complex induced Ginibre ensemble from the work of B. Rider [20]. In this reference, the author details an analytical method to determine the limiting distribution of the scaled spectral radius for matrices from the complex and symplectic Ginibre ensembles. A similar investigation is undertaken here for the scaled spectral radius and scaled minimum modulus of eigenvalues of matrices from the complex induced Ginibre ensemble. The analysis is conducted at the outer and inner edges of the ring (the eigenvalues support on the complex plane) defined with a rectangularity index assumed to be proportional to $N$, i.e., $L=\alpha N$, $\forall\alpha>0$. An exact fit between the empirical distribution and analytical formulation of these extreme moduli is acknowledged numerically. Additionally, the independence of the spectral radius and minimum modulus is studied in the limit as $N$ goes to infinity for these two non-Hermitian ensembles.
@@
Results stating the right and left tail eigenvalues distribution functions asymptotics for the complex elliptic Ginibre ensemble in the limit of weak non-Hermiticity is presented in [4]. The limiting tail distribution functions asymptotics of the eigenvalues modulus is also of interest. The similitude of the minimum modulus limiting (left and right) tail distributions between the two investigated non-Hermitian ensembles is explored in the present paper. This is the purpose of Section 4. Conclusions are set out in Section 5.
2. Limit theorems for the complex Ginibre ensemble
@@
Let $A$ denote a $N\times N$ complex Ginibre matrix whose symmetrised joint probability density function of its eigenvalues, $P_{N}(z_{1},\cdots,z_{N})$, is defined as in Section 1. The probability that the minimum of the eigenvalue moduli, here denoted $r_{min}^{(N)}(A)$, is greater than the radius $r\in\mathbb{R}^{+}$ is expressed as follows
$P\left(r_{min}^{(N)}(A)\geq r\right)=\int_{|z_{1}|\geq r}\cdots\int_{|z_{N}|\geq r}P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d^{2}z_{k}$
This probability corresponds to the probability that no eigenvalue lies inside the disk of radius $r$. It is a function of the radius $r$ and is initially defined analytically in reference from P. J. Forrester [6] with the following formula,
$P\left(r_{min}^{(N)}(A)\geq r\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ (10)
This result is presented in the reference from [2] whose authors applied the Gram’s formula [18] to get the probability, that no eigenvalue lies in the disk of radius $r$ centred at the origin, as a finite product of regularised upper incomplete Gamma functions. Equation (10) is derived with another approach in the present paper as follows. More precisely, the Andreief’s integration formula [7] is used to express analytically the gap probabilities $P(r_{min}^{(N)}(A)\geq a)$, $\forall a\in\mathbb{R}^{+*}$.
More precisely, from the definition of the eigenvalues joint probability density function defined in Section 1,
$P\left(r_{min}^{(N)}(A)\geq a\right)=\frac{1}{\Gamma_{N}}\int_{|z_{1}|>a}\cdots\int_{|z_{N}|>a}\det\left[z_{k}^{N-j}\right]_{j,k=1}^{N}\det\left[\bar{z}_{k}^{N-j}\right]_{j,k=1}^{N}\prod_{k=1}^{N}dm(z_{k})$
@@
Applying the Andreief’s integration formula,
@@
It is elementary to derive the probability density function $p_{r_{min}^{(N)}(A)}$ of the smallest modulus $r_{min}^{(N)}(A)$ from Equation (10). It is convenient to take the logarithm of the probability $P(r_{min}^{(N)}(A)\geq r)$ and find its first derivative with respect to $r$ to finally get the probability density function of the smallest modulus $r_{min}^{(N)}(A)$ as
@@
Using a similar approach as for the case of no scaling, the survival distribution function of scaled minimum of eigenvalue moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ for $N\times N$ complex Ginibre matrices $A$ has same analytical expression as for $r_{min}^{(N)}(A)$ replacing $r$ with $\sqrt{N}r$. The corresponding probability density function is the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ multiply by $\sqrt{N}$ at the point $\sqrt{N}r$, i.e.,
@@
The spectral radius of the real Ginibre matrices is investigated in reference [9]. It is experimentally stated that its value converges almost surely to the standard deviation of the i.i.d. entries of scaled matrices from this random matrix ensemble. The distribution of the spectral radius for a fixed size of complex Ginibre matrices is explored in the following.
@@
The cumulative distribution function of the spectral radius of eigenvalues, denoted $r_{max}^{(N)}(A)$, for the complex Ginibre ensemble is also easily retrieved with the use of the Andreief’s integration formula applied to the following multiple integrals equation
@@
The $N$-th gap probability, i.e., the probability that all the eigenvalues of a $N\times N$ complex Ginibre matrix lies in the disk of radius $r$, is the joint distribution of independent random variables $\gamma_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ [1], with $\alpha_{k}=k$ and $\delta_{k}=1$.
@@
where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in the work of E. Akarawak, I. Adeleke and R. Okafor [1].
The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (13).
Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density funvtion of the spectral radius $r_{max}^{(N)}(A)$ is,
@@
The scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is now considered in the following. The probability distribution function $P\left(\frac{r_{max}^{(N)}(A)}{\sqrt{N}}\leq r\right)$ of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is also derived using the Andreief’s integration formula [7],
where the random variable $Z^{(k)}=\sum_{j=1}^{k+1}Z_{j}$ follows a Gamma distribution with shape parameter $k+1$ and rate parameter $1$. The random variables $Z_{j}$, with $j\in\{1,...,k+1\}$, are independent and identically distributed. Each of the $Z_{j}$ follows a standard exponential distribution.
This is a known result presented in B. Rider’s work [20] from which he established the nature of the limiting distribution of the scaled spectral radius for matrices from the complex Ginibre ensemble in the limit as $N$ goes to infinity. The distribution of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for matrices from the complex Ginibre ensemble is the standard Gumbel distribution in the limit as $N$ goes to infinity. The limiting distribution of scaled $k$-th modulus has also been established for this random matrix ensemble as well as for the symplectic Ginibre ensemble in [21], in light of the framework presented in [20].

### v4

@@ Abstract.
The complex Ginibre ensemble and its generalisation, the complex induced Ginibre ensemble, have been well studied in the field of Random Matrix Theory. However, not all questions have formal responses. In the present work, asymptotic distributions and the independence of the modulus of the extreme eigenvalues are studied in the limit of large dimensions of random matrices for these non-Hermitian ensembles. They are derived with the use of Andreief’s integration formula and the known methodological approach defined for the study of the limiting distribution of the scaled spectral radius at the edge of the complex Ginibre ensemble. The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution. In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively. The limiting left tail distribution of the minimum modulus is the same for these matrix ensembles, with a rectangularity index of the complex induced Ginibre ensemble equal to zero. This phenomenon is also verified for the right tail distribution of this random minimum. The independence of the extreme moduli is formally established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes. This study extends knowledge in the field of Random Matrix Theory for random variables, like the minimum modulus of matrices from non-Hermitian ensembles, whose limiting stochastic dynamic has never been explored.
@@ 1. Introduction
The attractiveness of the Random Matrix Theory (RMT) lies in the possibility of using it as a means to model problems in high dimensions and perform related calculations analytically. This is mainly due to the invariance property of the probability distribution of certain matrix ensembles. As stated in [21], the statistical properties of the spectrum of random matrix ensembles are independent of the nature of the probability distribution that defines these ensembles in the limit of large sizes of these matrices. They only depend on the invariance of these distributions. Matrix ensembles that are characterised by invariant probability distribution are those from classical compact matrix groups studied in references [30] and [19]. The complex Ginibre ensemble is one of them and is a special case of the Ginibre-Girko ensemble with maximal non-Hermiticity as presented in [9] and [2]. Like the real and the symplectic Ginibre ensembles, this ensemble was introduced by J. Ginibre [11]. It was first defined as a mathematical concept to model phenomena from the physics of particles. Statistical properties of eigenvalues of matrices from this ensemble are studied to provide an understanding by analogy of the dynamics of nuclei. The distribution of eigenvalues of a complex Ginibre matrix is comparable to that of the distribution of the position of charges of a two-dimensional Coulomb gas in a harmonic oscillator potential, at a specific temperature corresponding to the Dyson index $\beta=2$.
The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$, whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density function
@@
where $(x_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$ and $(y_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$. The real and imaginary parts of the entries denoted $x_{ij}$ and $y_{ij}$ are independent random variables, each following a real Gaussian distribution $\mathcal{N}_{\mathbb{R}}(0,\frac{1}{2})$.
The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ the joint probability density function of the entries $J_{ij}$ is
@@
The term $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [17].
The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in closed form as
@@
where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of the eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$. The term $C_{N}$ is the normalisation factor defined as
@@
A topic which has been widely investigated in RMT is the limiting distribution of the eigenvalues of random matrix ensembles as $N$ goes to infinity. The notion of universality in this field has found relevance in establishing analytical expressions of the limiting distribution of random matrix eigenvalues. It is a mathematical concept asserting that the limiting distribution of the eigenvalues should not depend on the particular distribution of the random matrix entries [25]. This has led to several important results among which is referenced the circular law conjecture for non-Hermitian ensembles (cf. [3, 12, 25]).
Studies of the limiting distribution of the largest and smallest eigenvalues for Gaussian ensembles are presented in [27, 28, 29]. The distributions of eigenvalues (and their spacings) have been investigated with the computation of gap probabilities with respect to radial ordering for non-Hermitian random matrices and their chiral counterparts in [2]. The statistical properties of the spectral radius of matrices from non-Hermitian ensembles have also been studied in the literature, as it is for the complex Ginibre ensemble [22, 23] and the real Ginibre ensemble [24].
The present paper is structured as follows. Section 2 detailed the methodology used. The latter encompasses the application of Andreief’s integration formula [8] and the framework established by B. Rider in [22]. The results are presented in Section 3. Section 3.1 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble. It is especially performed for the spectral radius and the minimum modulus. The limiting left and right tail distributions of the minimum modulus are studied for large matrices in Section 3.1.1. Pursuing the analysis for this random matrix ensemble in the scaling limit $\sqrt{N}$ as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 3.1.2. The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process. The joint probability density of their respective modulus has been investigated in [18]. This result is extended in [16]. Employing the method introduced in [22] from which limit theorems are derived at the edge of the spectrum of the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [6] are presented in Section 3.2. This random matrix ensemble might find relevance for questions raised in different fields of physics. It would also define an ideal modelling framework for the expansion of the research related to non-Hermitian matrices. The complex induced Ginibre ensemble is a special case of the Feinberg-Zee ensemble and a generalisation of the complex Ginibre ensemble. The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is of the form
@@
The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V$ expressed as $V(y)=-\frac{\beta}{2}(y-L\log(y))$, where $L$ is a non-negative parameter called the rectangularity index, and $\beta=1$ or $\beta=2$ for the real or complex induced Ginibre ensembles, respectively.
@@
$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)$ (6)
where $\beta=2$ and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$. The factor $\Gamma^{(\beta)}$ is the normalisation constant
@@
In the limit of large matrix dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble, thus referenced as the circular law and presented in the work from V. L. Girko [12] and the paper from T. Tao, V. Vu and M. Krishnapur [25]. The joint probability density function of the eigenvalues of the complex induced Ginibre matrix $G$ is
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)$ (8)
@@
It is understood from the definition of the joint probability density function of the eigenvalues that there is a repulsion of the eigenvalues from the origin of the complex plane due to the term $\det(GG^{\dagger})^{L}=\prod_{j=1}^{N}|\lambda|^{2L}$ appearing in the formula. This phenomenon is explained from a mathematical perspective with the formulation of the mean eigenvalue density, i.e., $\langle\rho_{N}(\lambda)\rangle=R_{1}(\lambda)$ where $R_{1}(\lambda)$ is the one-point correlation function of the eigenvalues. More precisely taking its limit as $N$ goes to infinity and for a non-negative rectangularity index $L$, the distribution of the eigenvalues is uniform and supported by a ring centred at the origin of the complex plane with an outer circle of radius $r_{out}=\sqrt{L+N}$ and an inner circle of radius $r_{in}=\sqrt{L}$. This phenomenon refers to the ”Single Ring Theorem” stating that, in the limit of large matrix dimensions, the empirical distribution of the eigenvalues has for support a single ring [5] [15]. For small rectangularity indexes $L$ (close to zero), the density of the eigenvalues is uniform over the disk of radius $\sqrt{N}$ with a mean eigenvalue density converging to the mean eigenvalue density of the complex Ginibre ensemble.
@@
A study is performed Section 3.2 for the complex induced Ginibre ensemble from the work of B. Rider [22]. In this reference, the author details an analytical method to determine the limiting distribution of the scaled spectral radius for matrices from the complex and symplectic Ginibre ensembles. A similar analysis is undertaken in Section 3.2.1 for the scaled spectral radius and scaled minimum modulus of eigenvalues of matrices from the complex induced Ginibre ensemble. The study is conducted at the outer and inner edges of the ring (the eigenvalues support on the complex plane) defined with a rectangularity index assumed to be proportional to $N$, i.e., $L=\alpha N$, $\forall\alpha>0$. An exact fit between the empirical distribution and analytical formulation of these extreme moduli is acknowledged numerically. Additionally, the independence of the spectral radius and minimum modulus is studied in the limit as $N$ goes to infinity for these two non-Hermitian ensembles in Section 3.2.2.
@@
The limiting tail distribution functions of the eigenvalues modulus is also of interest in the field. Results stating the right and left tail eigenvalues distribution function asymptotics for the complex elliptic Ginibre ensemble in the limit of weak non-Hermiticity is presented in [4]. The similitude of the minimum modulus limiting right tail distributions between the two non-Hermitian ensembles at scrutiny is explored in Section 3.3 of the present paper. The same analysis is performed for the left tail distribution of this extreme modulus. A discussion and a list of conclusions are set out in Section 4.
2. Methods
Let $P\left(r_{min}^{(N)}(A)\geq r\right)$ denote the survival distribution function of the minimum modulus $r_{min}^{(N)}(A)$ and
@@ 2. Methods
$P\left(r_{max}^{(N)}(A)\leq r\right)$ the cumulative distribution function of the spectral radius $r_{max}^{(N)}(A)$, for a $N\times N$ complex matrix $A$. These distribution functions are derived analytically, in the present paper, with the use of Andreief’s integration formula presented in [8]. A straightforward application is as follows. Let $A$ denote a $N\times N$ complex Ginibre matrix whose the symmetrised joint probability density function of its eigenvalues, denoted $P_{N}(z_{1},\cdots,z_{N})$, is defined as in Section 1 with Equation (3). The probability that the minimum of the eigenvalue moduli, here denoted $r_{min}^{(N)}(A)$, is greater than the radius $a\in\mathbb{R}^{+}$, is
$P\left(r_{min}^{(N)}(A)\geq a\right)$ $=\int_{|z_{1}|\geq a}\cdots\int_{|z_{N}|\geq a}P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d^{2}z_{k}$
$=\frac{1}{\Gamma_{N}}\int_{|z_{1}|>a}\cdots\int_{|z_{N}|>a}\det\left[z_{k}^{N-j}\right]_{j,k=1}^{N}\det\left[\bar{z}_{k}^{N-j}\right]_{j,k=1}^{N}\prod_{k=1}^{N}dm(z_{k})$
@@
Then, performing a change of variables in polar coordinates gives,
@@
This probability corresponds to the probability that no eigenvalue lies inside the disk of radius $a$ and is analytically formulated as
$P\left(r_{min}^{(N)}(A)\geq a\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,a^{2})}{\Gamma(k+1)}$ (10)
This result is initially stated without proof and without scaling applied to the minimum modulus in the reference [14], and for which asymptotic expansions are studied in [7]. It is also presented in the reference [2] in which the Gram’s formula [20] is applied to get this probability (the probability that no eigenvalue lies in the disk of radius $a$ centred at the origin of the complex plane) as a finite product of regularised upper incomplete Gamma functions. It is straightforward to derive the probability density function $p_{r_{min}^{(N)}(A)}$ of the minimum modulus $r_{min}^{(N)}(A)$ from Equation (10). More precisely, it is convenient to take the logarithm of the probability $P(r_{min}^{(N)}(A)\geq r)$ and find its first derivative with respect to $r$. Let the survival distribution function $P(r_{min}^{(N)}(A)\geq r)$ denote the function $H(r)$ of the radius $r$. The corresponding probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is $p_{r_{min}^{(N)}(A)}(r)=-\frac{\text{d}H(r)}{\text{d}r}$.
Furthermore,
$\frac{\text{d}H(r)}{\text{d}r}$ $=H(r)\frac{\text{d}}{\text{d}r}\left(\log\left[H(r)\right]\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\left[\sum_{k=0}^{N-1}\frac{\partial}{\partial r}\left[\log\left(\Gamma(k+1,r^{2})\right)\right]\right]$
with,
$\frac{\partial\log\left[\Gamma(k+1,r^{2})\right]}{\partial r}$ $=\frac{1}{\Gamma(k+1,r^{2})}\frac{\partial\Gamma(k+1,r^{2})}{\partial r}$
It is known that $\frac{\partial\Gamma(s,x)}{\partial x}=-x^{s-1}e^{-x}$.
Thus,
$\frac{\text{d}H(r)}{\text{d}r}$ $=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}\frac{\partial\Gamma(j+1,r^{2})}{\partial r}\right]$
where
$\frac{\partial\Gamma(j+1,r^{2})}{\partial r}=-2r^{2j+1}e^{-r^{2}}$
Finally, the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is,
@@
Using a similar approach as for the case of no scaling, the survival distribution function of scaled minimum of eigenvalue moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ for an $N\times N$ complex Ginibre matrix $A$ has same analytical expression as for $r_{min}^{(N)}(A)$ replacing $r$ with $\sqrt{N}r$. The corresponding probability density function is the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ multiply by $\sqrt{N}$ at the point $\sqrt{N}r$, i.e.,
@@
The spectral radius of matrices from the real Ginibre ensemble is investigated in reference [10]. It is experimentally stated that its value converges almost surely to the standard deviation of the i.i.d. entries of scaled matrices from this random matrix ensemble. The distribution function of the spectral radius for a fixed size of complex Ginibre matrices is explored in the following.
@@
Let $A$ denote an $N\times N$ complex Ginibre matrix. Let $r_{max}(A)^{(N)}$ denote the spectral radius.
Then,
@@
The cumulative distribution function of the spectral radius $r_{max}^{(N)}(A)$ for the complex Ginibre ensemble is also easily retrieved with the use of Andreief’s integration formula applied to the following multiple integrals equation
@@
The $N$-th gap probability, i.e., the probability that all the eigenvalues of an $N\times N$ complex Ginibre matrix lies in the disk of radius $r$, is the joint distribution of independent random variables $\gamma_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ [1], with $\alpha_{k}=k$ and $\delta_{k}=1$.
@@
where the random variables $\gamma_{k}$ are independent and each follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$, with parameters $\alpha_{k}=k$ and $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in the work from E. Akarawak, I. Adeleke and R. Okafor [1]. ∎
The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (13). Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density function of the spectral radius $r_{max}^{(N)}(A)$ is,